Use mathematical induction to prove that 2n > n2 , for n > 5 .
1
Expert's answer
2021-10-13T17:23:53-0400

Using mathematical induction let us prove that "2^n > n^2 ," for "n > 5 ."


Let "n=6." Then "2^6=64>36=6^2," and hence for "n=6" the statement is true.


Suppose that the statement is true for "n=k>5," that is "2^k > k^2."


Let us prove this statement for "n=k+1."


For this firstly let us show that "k^2>2k+1" for "k>5." In this case "k-1>4," and hence "(k-1)^2>16." It follows that "k^2-2k+1>16," and thus "k^2-2k>15>1." Therefore, "k^2>2k+1."


Taking into acount that "2^k > k^2," we conclude that

"2^{k+1}=2\cdot2^k>2k^2=k^2+k^2>k^2+2k+1=(k+1)^2,"

and the statement is true for "n=k+1."


We conclue that the statement "2^n > n^2" is true for all natural numbers "n > 5 ."